[Brahms-l] Shift report 20050615 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Jun 15 2005 - 15:59:07 EDT
Shifter leader was Theresa Wilkerson
Shifters were 


1. Machine status normal during shift.<br>
2. Run #15496 & 15497 pedestal runs. Run #15498-15508 data runs of varying usefulness. Refer to E-log notes.<br>
3. Power supplies for D1 and D2 magnets being worked on for much of the day. Took as much data as possible, but the FFS data is corrupt in later runs. Refer to E-log.<br>

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Jun 15 21:23:20 2005

This archive was generated by hypermail 2.1.8 : Wed Jun 15 2005 - 21:23:32 EDT